In a pre-2009 business combination, Acme Company acquired all of Brem Company’s assets and liabilities for cash. After the combination Acme formally dissolved Brem. At the acquisition date, the following book and fair values were available for the Brem Company accounts:

Book Values Fair Values
Current assets $ 81,800 $ 81,800
Equipment 131,000 198,000
Trademark 0 352,000
Liabilities (67,800 ) (67,800 )
Common stock (100,000 )
Retained earnings (45,000 )

In addition, Acme paid an investment bank $31,200 cash for assistance in arranging the combination.

  1. Using the legacy purchase method for pre-2009 business combinations, prepare Acme’s entry to record its acquisition of Brem in its accounting records assuming the following cash amounts of $690,400 and $439,400 were paid to the former owners of Brem.
  2. How would these journal entries change if the acquisition occurred post-2009 and therefore Acme applied the acquisition method?

Journal entry

  • 1

    Record the acquisition of Brem using the purchase method assuming $690,400 was paid to the former owners of Brem

  • 2

    Record the acquisition of Brem using the purchase method assuming $439,400 was paid to the former owners of Brem

  • 3

    Record the acquisition of Brem using the acquisition method assuming $690,400 was paid to the former owners of Brem

  • 4

    Record the expenses related to the combination using the acquisition method assuming $690,400 was paid to the former owners of Brem

  • 5

    Record the acquisition of Brem using the acquisition method assuming $439,400 was paid to the former owners of Brem

  • 6

  • Record the expenses related to the combination using the acquisition method assuming $439,400 was paid to the former owners of Brem
Answer #1

Answer Page No 9 @ sehen se cash payment period to enter the scenes bo ksiloyDO Goodwill = ( cash amount + acquisition coola)Bargain purchase Page. No @ = ( 439400 + 31200) -(81800 + 198000 + 352000 – 67800 00 600 - 664000 = - $93400 Trademarila = 35Credit Pageino eshen the cast amount paid to foren coners id $990400 Goodwill - Consideration transferred – (fair value of nePage. No 4 = 4 39400 - (81800 + 198000 + 352000 - 67800) 439400 - 564000 = = - $ (24600 Credit General journal Current asselo